# Collaborative and parametric insurance on the Ethereum blockchain

### v2

@@ 3.2 Parametric insurance portfolio over time
A parametric insurance portfolio is a collection of parametric insurance contracts $\{(S_{i},T_{i},Q_{i},\overline{Q}_{i},l_{i})\}_{i\geq 1}$. These contracts may be associated with their own observable quantities $Q_{i}:=\left(Q_{t}^{i}\right)_{t\geq 0}$ defined on a common probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$. Let us take a concrete situation in 2 by following up on 1.
@@
In addition to Strasbourg, we wish to offer coverage against rain episodes in the city of Marseille. We assume that Strasbourg and Marseille are sufficiently far apart (see the map on Figure 5) so that the precipitation levels of these cities are not correlated.
@@
Each contract $i$ is then associated with a location so that $Q^{i}\in\{Q^{\text{Marseille}},Q^{\text{Strasbourg}}\}$ and a date $T_{i}$ for all $i\geq 1$. We assume that $Q^{\text{Marseille}}$ and $Q^{\text{Strasbourg}}$ are mutually independent. We assume that $S_{i}<S_{i}^{\ast}$ for each contract so as to model the precipitation height via a compound Poisson-gamma distribution at each location. The models are fitted to the same data as in 1 and result in two different risk profiles as illustrated on Figure 6, where we plot the probability $\mathbb{P}(Q_{T}>5)$ as a function of the month of $T$ for each location.
@@
It corresponds to the future losses given the active policies in our portfolio at time $t$. The premiums collected to compensate for this liability are given by
@@ 3.3.1 Model points creation to handle dependency
The main challenge associated with calculating the risk capitals (3) lies in the study of the distribution of the liability $(L_{t})_{t\geq 0}$, more specifically the likely dependency among the Bernoulli variables. We show in 3 how to gather two correlated contracts into one within the framework of 1 and 2.
@@
Consider two contracts $\left(S_{i_{1}},T_{i_{1}},Q_{i_{1}},\overline{Q}_{i_{1}},l_{i_{1}}\right)$ and $\left(S_{i_{2}},T_{i_{2}},Q_{i_{2}},\overline{Q}_{i_{2}},l_{i_{2}}\right)$ associated with the same location $Q=Q_{i_{1}}=Q_{i_{2}}$ and the same event date $T=T_{i_{1}}=T_{i_{2}}$. The two contracts do not have the same threshold nor the same underwriting time. We assume without loss of generality that $\overline{Q}_{i_{1}}\leq\overline{Q}_{i_{2}}$ and $S_{i_{1}}<S_{i_{2}}$. Denote by
@@
Following up on 3, for $t\geq 0$, we partition our insurance portfolio $\left\{\left(S_{i},T_{i},Q_{i},\overline{Q}_{i},l_{i}\right)\right\}_{i=1,\ldots,N_{t}}$ into $M_{t}$ groups $\text{MP}_{1,t},\ldots,\text{MP}_{M_{t},t}$ of correlated contracts. Each class is associated with a payout $Z_{j,t}$, $j=1,\ldots,M_{t}$. The process $(M_{t})_{t\geq 0}$ counts the number of model points defined up to time $t\geq 0$. The liability $(L_{t})_{t\geq 0}$ is rewritten as
@@
The liability is now a sum of independent, discrete, and positive random variables. The concept of model points has been introduced in the European directive Solvency II to speed up the calculations of Best Estimate Liabilities associated with life insurance portfolios when using cash-flow projection models. Grouping and aggregating insurance contracts has become a common practice among actuaries, and various methods have been documented in the actuarial science literature, see for instance the works of Goffard and Guerrault [16], Blanchet-Scalliet et al. [5], Kiermayer and Weiß [19], Gweon et al. [18], Gweon and Li [17]. Here, the creation of model points is a workaround for the dependency of the Bernoulli variables, which allows us to study the distribution of the liability in a tractable way as shown in Section 3.3.2.
@@ 3.3.2 Recursive calculation of the solvency capitals
Provided that $M_{t}^{\text{active}}=n$, the payouts of $\text{MP}_{1},\ldots,\text{MP}_{n}$, at time $t\geq 0$ are denoted by $Z_{1},\ldots,Z_{n}$. Consequently, the distribution of the liability $L_{t}=\sum_{k=1}^{n}Z_{k}$ is represented as the sum of independent random variables. While it is possible to compute the exact distribution using combinatorial analysis, this method becomes impractical as $n$ increases.
A more efficient approach involves using a Fast Fourier Transform (FFT) procedure. After obtaining the probability distribution of $L_{t}$, we can compute its cumulative distribution function and inverse it using a bisection algorithm. However, despite its effectiveness, FFT cannot be implemented on the blockchain due to the limitations of smart contract functions, which are unable to encode complex operations. This constraint necessitates performing solvency capital calculations off-chain, requiring intervention from the smart contract manager. Furthermore, any changes in the portfolio at a later time $s\geq t$ would require recomputing the probability distribution of the liability $L_{s}$ to determine the Solvency Capital Requirement ($\text{SCR}_{s}$) or Minimum Capital Requirement ($\text{MCR}_{s}$).
@@
Suppose the sequence of random variables $(Z_{k})_{k=1,\ldots,n}$ satisfies the Lindeberg-Feller condition:
$\lim_{n\to\infty}\frac{1}{s_{n}^{2}}\sum_{k=1}^{n}\mathbb{E}\left[(Z_{k}-\mu_{k})^{2}\mathbb{I}_{|Z_{k}-\mu_{k}|>\epsilon s_{n}}\right]=0,\quad\text{for every }\epsilon>0.$
Then, we have the following convergence in distribution:
@@
where $\mathsf{Normal}(0,1)$ represents the standard normal distribution and $\overset{\mathcal{D}}{\longrightarrow}$ denotes convergence in distribution.
Given that the random variables $Z_{k}$ are bounded, it is straightforward to verify that the Lyapunov condition, which is stronger than the Lindeberg-Feller condition, holds. The Lyapunov condition is satisfied if there exists a $\delta>0$ such that:
We note that
@@
respectively, where $\gamma_{1}$ and $\gamma_{2}$ denote the skewness and excess kurtosis of $L_{t}$, respectively. Note that the approximation in (5) is referred to as the second-order Cornish-Fisher approximation. We then decide on a threshold $\bar{n}$ such that if $n>\bar{n}$, we compute the solvency capital using the approximations in (5) or (6). Otherwise, we fall back to:
@@
The determination of such a threshold can be based on a brief simulation study, as illustrated in 4.
@@
The first five rows of such a synthetic portfolio are provided in Table 2.
@@
We simulate 100 synthetic portfolios and calculate the relative errors between the exact and approximated solvency capitals for an increasing number of model points, $n$. The order of the quantiles are set to $q_{\alpha_{\text{MCR}}}=0.85$ and $q_{\alpha_{\text{SCR}}}=0.995$. The results of these simulations are displayed in Figure 7.
@@
The normal approximation becomes acceptable for the Minimum Capital Requirement (MCR) relatively quickly, but the relative error for the Solvency Capital Requirement (SCR) converges more slowly. Overall, adjusting for skewness is beneficial for estimating both the MCR and SCR. The level of error falls below 5% when $n\geq 15$ and drops below 1% for $n\geq 30$. If we consider a 5% error level to be acceptable, then the third-order Cornish-Fisher approximation could be used as soon as the number of model points reaches 15.
@@
It is common practice among insurers to limit the underwriting of highly correlated risks. This can be easily implemented by preventing underwriting if the contract is aggregated within a model point that already contains too many contracts. Note that we implicitly do this in 4, as the number of correlated risks is limited to 10. Additionally, the condition could be based on the maximum possible loss associated with the model point.
The conclusions and resulting recommendations of 4 are valid for the specific case under consideration. A similar study could be conducted to adapt the threshold for the number of model points and guide the selection of an appropriate order for the Cornish-Fisher approximation for another situation.
@@ 4 Smart contract description
A smart contract is a system whose state evolves according to the actions of blockchain users. The roles of the users of our smart contract are outlined in Section 4.1. The stochastic processes used to monitor the state of the system are defined in Section 4.2. We introduce the methods that allow users to interact with the smart contract in Section 4.3.
@@ 4.1.1 Smart contract owner
The smart contract owner is responsible for designing the insurance protocol by writing the Solidity code for the smart contract. While deploying the smart contract on the Ethereum blockchain, the owner initializes three parameters: the premium loading $\eta$, the risk levels $\alpha_{\text{SCR}}$ and $\alpha_{\text{MCR}}$ associated with the solvency capitals, as defined in Section 3, and the lower bound $m$ for the number of active model points to use the normal approximation of Section 3.3 to calculate the solvency capitals. Additionally, the contract owner acts as the claim assessor or data oracle by inputting the observed quantity at the time of the event—in this case, the precipitation height. In the blockchain ecosystem, an oracle refers to an off-chain data provider that a smart contract can query. This feature is not implemented in our current solution, it would be beneficial to do so in order to further automate our workflow.
@@ 4.1.2 Surplus providers
Surplus providers, also known as investors, contribute financially to the smart contract to constitute the solvency capital. These contributions are denominated in ETH, the native cryptocurrency of the Ethereum blockchain. In return for their investment, investors receive tokens that represent their proportional share of the surplus. Surplus providers have the option to withdraw a portion of their funds by burning these tokens in exchange for an equivalent amount of ETH, with the exchange rate defined in the Section 4.2. Withdrawals are permitted provided that the funds remaining in the smart contract do not fall below a specified threshold, known as the Solvency Capital Requirement (SCR).
@@ 4.1.3 Policyholders
An insurance agreement is only concluded if the Solvency Capital Requirement (SCR), after accounting for the new risk, does not exceed the funds currently locked in the smart contract. Should the settlement of a claim lead to the smart contract’s funds falling below the Minimum Capital Requirement (MCR), the system enters a state of bankruptcy. In such instances, priority is given to reimbursing policyholders with the remaining funds. Any funds that remain after all policyholder reimbursements are distributed to investors, proportional to their token holdings.
@@ 4.2 State variables
The state variables of the smart contract include the stochastic processes defined in Section 3, supplemented by additional processes defined on the filtered probability space $(\Omega,\mathcal{F},(\mathcal{F})_{t\geq 0},\mathbb{P})$.
@@
Note that $B_{t}\geq X_{t}$ for every $t\geq 0$. The surplus $(X_{t})_{t\geq 0}$ is compared against $\text{SCR}_{t}$ and $\text{MCR}_{t}$. If $X_{t}\leq\text{SCR}_{t}$, no new policies may be underwritten, and investors are prohibited from withdrawing funds. If $X_{t}\leq\text{MCR}_{t}$, the smart contract ceases operations, and the remaining balance is redistributed—first to policyholders and then to investors, as detailed in Section 4.3.4. Active insurance policies at the time of termination are cancelled.
@@ 4.3 Method of the smart contract
The state of a smart contract changes according to the actions of blockchain users. In our parametric insurance example, these users include the contract owner, surplus providers, and policyholders (see Section 4.1). Participants interact with the smart contract through specific functions, known as methods, which modify the system’s state. The remainder of this section describes each possible action and its impact on the stochastic processes introduced thus far.
@@ 4.3.2 Underwriting a parametric insurance policy
where $\theta=\mathbb{P}(Q_{T}>\overline{Q}|\mathcal{F}_{S})$. The contract can only be underwritten if the surplus of the smart contract is sufficient to cover the resulting increase in the Solvency Capital Requirement (SCR).
@@
$\widetilde{M}^{\text{active}}_{t}=\begin{cases}M^{\text{active}}_{t},&\text{ifthe new contract is grouped into an active model point},\\ M^{\text{active}}_{t}+1,&\text{otherwise}.\end{cases}$
@@ 4.3.4 Contract resolution and claim settlement
Recall that $S_{i}$ is the time at which contract $i$ was underwritten. The number of active model poinst is updated as follows:
$\widetilde{M}^{\text{active}}_{t}=\begin{cases}M^{\text{active}}_{t}-1,&\text{if the contract was the last contract of a given model point},\\ M^{\text{active}}_{t},&\text{otherwise}.\end{cases}$
The Solvency Capital Requirement (SCR) and Minimum Capital Requirement (MCR) are also updated as:
@@
Now suppose that the surplus downcrosses the MCR as $\widetilde{X}_{t}\leq\widetilde{\text{MCR}}_{t}$. If this occurs, the smart contract is reset. The outstanding balance is first used to reimburse the policyholders of active contracts whose status becomes "cancelled". The total amount is therefore given by:

### v3

@@ 3.2 Parametric insurance portfolio over time
A parametric insurance portfolio is a collection of parametric insurance contracts $\{(S_{i},T_{i},Q_{i},\overline{Q}_{i},l_{i})\}_{i\geq 1}$. These contracts may be associated with their own observable quantities $Q_{i}:=\left(Q_{t}^{i}\right)_{t\geq 0}$ defined on a common probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$. Let us take a concrete situation in ˜2 by following up on ˜1.
@@
In addition to Strasbourg, we wish to offer coverage against rain episodes in the city of Marseille. We assume that Strasbourg and Marseille are sufficiently far apart (see the map on Figure˜5) so that the precipitation levels of these cities are not correlated.
@@
Each contract $i$ is then associated with a location so that $Q^{i}\in\{Q^{\text{Marseille}},Q^{\text{Strasbourg}}\}$ and a date $T_{i}$ for all $i\geq 1$. We assume that $Q^{\text{Marseille}}$ and $Q^{\text{Strasbourg}}$ are mutually independent. We assume that $S_{i}<S_{i}^{\ast}$ for each contract so as to model the precipitation height via a compound Poisson-gamma distribution at each location. The models are fitted to the same data as in ˜1 and result in two different risk profiles as illustrated on Figure˜6, where we plot the probability $\mathbb{P}(Q_{T}^{i}>5)$ as a function of the month of $T$ for each location $i$.
@@
It corresponds to the future losses given the active policies in our portfolio at time $t$.
Remark 3.6.
Note that the assumption of fixed-parameter Bernoulli variables $Y_{i}$ with parameters is an approximation. In reality, the event probabilities follow stochastic processes, recall the definition (2). A more precise approach would involve dynamically updating these probabilities when analyzing the liability distribution at time $t$. However, we adopt this simplification to avoid the prohibitive transaction costs that would arise from real-time updates in a blockchain environment, which would require frequent oracle queries and solvency capital reevaluations.
The premiums collected to compensate for this liability are given by
@@ 3.3.1 Model points creation to handle dependency
The main challenge associated with calculating the risk capitals (4) lies in the study of the distribution of the liability $(L_{t})_{t\geq 0}$, more specifically the likely dependency among the Bernoulli variables. We show in ˜3 how to gather two correlated contracts into one within the framework of ˜1 and ˜2.
@@
Consider two contracts $\left(S_{i_{1}},T_{i_{1}},Q_{i_{1}},\overline{Q}_{i_{1}},l_{i_{1}}\right)$ and $\left(S_{i_{2}},T_{i_{2}},Q_{i_{2}},\overline{Q}_{i_{2}},l_{i_{2}}\right)$ associated with the same location $Q=Q_{i_{1}}=Q_{i_{2}}$ and the same event date $T=T_{i_{1}}=T_{i_{2}}$. The two contracts do not have the same threshold nor the same underwriting time. We assume without loss of generality that $\overline{Q}_{i_{1}}\leq\overline{Q}_{i_{2}}$ and $S_{i_{1}}<S_{i_{2}}$. We assume that $S_{i_{1}}$ and $S_{i_{2}}$ are both sufficiently far from the event date $T$. Denote by
@@
Following up on ˜3, for $t\geq 0$, we partition our insurance portfolio $\left\{\left(S_{i},T_{i},Q_{i},\overline{Q}_{i},l_{i}\right)\right\}_{i=1,\ldots,N_{t}}$ into $M_{t}$ groups $\text{MP}_{1,t},\ldots,\text{MP}_{M_{t},t}$ of correlated contracts. The classes are associated with independent payouts $Z_{j,t}$, $j=1,\ldots,M_{t}$. The process $(M_{t})_{t\geq 0}$ counts the number of model points defined up to time $t\geq 0$. The liability $(L_{t})_{t\geq 0}$ is rewritten as
@@
The liability is now a sum of independent, discrete, and positive random variables. The concept of model points has been introduced in the European directive Solvency II to speed up the calculations of Best Estimate Liabilities associated with life insurance portfolios when using cash-flow projection models. Grouping and aggregating insurance contracts has become a common practice among actuaries, and various methods have been documented in the actuarial science literature, see for instance the works of Goffard and Guerrault [16], Blanchet-Scalliet et al. [5], Kiermayer and Weiß [19], Gweon et al. [18], Gweon and Li [17]. Here, the creation of model points is a workaround for the dependency of the Bernoulli variables, which allows us to study the distribution of the liability in a tractable way as shown in Section˜3.3.2.
@@ 3.3.2 Recursive calculation of the solvency capitals
Provided that $M_{t}^{\text{active}}=n$, the payouts of $\text{MP}_{1},\ldots,\text{MP}_{n}$, at time $t\geq 0$ are denoted by $Z_{1},\ldots,Z_{n}$. Consequently, the distribution of the liability $L_{t}=\sum_{k=1}^{n}Z_{k}$ is represented as the sum of independent random variables. While it is possible to compute the exact distribution using combinatorial analysis, this method becomes impractical as $n$ increases. We use instead a Fast Fourier Transform (FFT) procedure. After obtaining the probability distribution of $L_{t}$, we can compute its cumulative distribution function and inverse it using a bisection algorithm. However, despite its effectiveness, FFT cannot be implemented on the blockchain due to the limitations of smart contract functions, which are unable to encode complex operations. This constraint necessitates performing solvency capital calculations off-chain, requiring intervention from the smart contract manager. Furthermore, any changes in the portfolio at a later time $s\geq t$ would require recomputing the probability distribution of the liability $L_{s}$ to determine the Solvency Capital Requirement ($\text{SCR}_{s}$) or Minimum Capital Requirement ($\text{MCR}_{s}$).
@@
Since the random variables $(Z_{k})_{k=1,\ldots,n}$ are bounded, they satisfy Lyapunov’s condition that reads as
Thus, the following convergence result holds:
@@
where $\mathsf{Normal}(0,1)$ represents the standard normal distribution and $\overset{\mathcal{D}}{\longrightarrow}$ denotes convergence in distribution. We note that
@@
respectively, where $\gamma_{1}$ and $\gamma_{2}$ denote the skewness and excess kurtosis of $L_{t}$. Note that the approximation in (6) is referred to as the second-order Cornish-Fisher approximation. We then decide on a threshold $\bar{n}$ such that if $n>\bar{n}$, we compute the solvency capital using the approximations in (6) or (7). Otherwise, we fall back to:
@@
The determination of such a threshold can be based on a brief simulation study, as illustrated in ˜4.
@@
The first five rows of such a synthetic portfolio are provided in Table˜2.
@@
We simulate 100 synthetic portfolios and calculate the relative errors between the exact and approximated solvency capitals for an increasing number of model points, $n$. The order of the quantiles are set to $q_{\alpha_{\text{MCR}}}=0.85$ and $q_{\alpha_{\text{SCR}}}=0.995$. The results of these simulations are displayed in Figure˜7.
@@
The normal approximation becomes acceptable for the MCR relatively quickly, but the relative error for the SCR converges more slowly. Overall, adjusting for skewness is beneficial for estimating both the MCR and SCR. The level of error falls below 5% when $n\geq 15$ and drops below 1% for $n\geq 30$. If we consider a 5% error level to be acceptable, then the third-order Cornish-Fisher approximation could be used as soon as the number of model points reaches 15.
@@
It is common practice among insurers to limit the underwriting of highly correlated risks. This can be easily implemented by preventing underwriting if the contract is aggregated within a model point that already contains too many contracts. Note that we implicitly do this in ˜4, as the number of correlated risks is limited to 10. Additionally, the condition could be based on the maximum possible loss associated with the model point.
The conclusions and resulting recommendations of ˜4 are valid for the specific case under consideration. A similar study could be conducted to adapt the threshold for the number of model points and guide the selection of an appropriate order for the Cornish-Fisher approximation for another situation.
@@ 4 Smart contract description
A smart contract is a system whose state evolves according to the actions of blockchain users. The roles of the users of our smart contract are outlined in Section˜4.1. The stochastic processes used to monitor the state of the system are defined in Section˜4.2. We introduce the methods that allow users to interact with the smart contract in Section˜4.3.
@@ 4.1.1 Smart contract owner
The smart contract owner is responsible for designing the insurance protocol by writing the Solidity code for the smart contract. While deploying the smart contract on the Ethereum blockchain, the owner initializes three parameters: the premium loading $\eta$, the risk levels $\alpha_{\text{SCR}}$ and $\alpha_{\text{MCR}}$ associated with the solvency capitals, as defined in Section˜3, and the lower bound $m$ for the number of active model points to use the normal approximation of Section˜3.3 to calculate the solvency capitals. Additionally, the contract owner acts as the claim assessor or data oracle by inputting the observed quantity at the time of the event—in this case, the precipitation height. In the blockchain ecosystem, an oracle refers to an off-chain data provider that a smart contract can query. This feature is not implemented in our current solution, it would be beneficial to do so in order to further automate our workflow.
@@ 4.1.2 Surplus providers
Surplus providers, also known as investors, contribute financially to the smart contract to constitute the solvency capital. These contributions are denominated in ETH, the native cryptocurrency of the Ethereum blockchain. In return for their investment, investors receive tokens that represent their proportional share of the surplus. Surplus providers have the option to withdraw a portion of their funds by burning these tokens in exchange for an equivalent amount of ETH, with the exchange rate defined in the Section˜4.2. Withdrawals are permitted provided that the funds remaining in the smart contract do not fall below a specified threshold, known as the Solvency Capital Requirement (SCR).
@@ 4.1.3 Policyholders
An insurance agreement is only concluded if the SCR, after accounting for the new risk, does not exceed the funds currently locked in the smart contract. Should the settlement of a claim lead to the smart contract’s funds falling below the MCR, the system enters a state of bankruptcy. In such instances, priority is given to reimbursing policyholders with the remaining funds. Any funds that remain after all policyholder reimbursements are distributed to investors, proportional to their token holdings.
@@ 4.2 State variables
The state variables of the smart contract include the stochastic processes defined in Section˜3, supplemented by additional processes defined on the filtered probability space $(\Omega,\mathcal{F},(\mathcal{F})_{t\geq 0},\mathbb{P})$.
@@
Note that $B_{t}\geq X_{t}$ for every $t\geq 0$. The surplus $(X_{t})_{t\geq 0}$ is compared against $\text{SCR}_{t}$ and $\text{MCR}_{t}$. If $X_{t}\leq\text{SCR}_{t}$, no new policies may be underwritten, and investors are prohibited from withdrawing funds. If $X_{t}\leq\text{MCR}_{t}$, the smart contract ceases operations, and the remaining balance is redistributed—first to policyholders and then to investors, as detailed in Section˜4.3.4. Active insurance policies at the time of termination are cancelled.
@@ 4.3 Method of the smart contract
The state of a smart contract changes according to the actions of blockchain users. In our parametric insurance example, these users include the contract owner, surplus providers, and policyholders (see Section˜4.1). Participants interact with the smart contract through specific functions, known as methods, which modify the system’s state. The remainder of this section describes each possible action and its impact on the stochastic processes introduced thus far.
@@ 4.3.2 Underwriting a parametric insurance policy
where $\theta=\mathbb{P}(Q_{T}>\overline{Q}|\mathcal{F}_{S})$. The contract can only be underwritten if the surplus of the smart contract is sufficient to cover the resulting increase in the SCR.
@@
[…]
@@ 4.3.4 Contract resolution and claim settlement
Recall that $S_{i}$ is the time at which contract $i$ was underwritten. The number of active model points is updated as follows:
[…]
The SCR and MCR are also updated as:
@@
Now suppose that the surplus downcrosses the MCR as $\widetilde{X}_{t}\leq\widetilde{\text{MCR}}_{t}$. If this occurs, the smart contract is reset. The outstanding balance is first used to reimburse the policyholders of active contracts whose status become "cancelled". The total amount is therefore given by: